If Ben invests $4000 at 4% interest per year, how much additional money must he invest at 5.5% annual interest to ensure that the interest he receive each year is 4.5 of the total amount invested?
Answer by checkley77(12569)

.055x+.04*4,000=.045(4,000+x)
.055x+160=180+.045x
.055x-.045x=180-160
.01x=20
x=20/.01
x=2,000 needs to be invested @ 5.5%.
Proof:
.055*2,000+.04*4,000=.045(4,000+2,000)
110+160=.045*6,000
270=270